OpenStudy (anonymous):

solve for x: sqrt{x+7}=x+5

OpenStudy (anonymous):

check your answer

OpenStudy (anonymous):

Squeaky Think About it !!! How WILL YOU REMOVE A SQRT(1/2).........

OpenStudy (anonymous):

Square both sides, dude. Then foil.

OpenStudy (anonymous):

sqrt{x+7}=x+5 [square both sides] x +7 = (x+5)^2 x +7 = x^2 + 10x +25 0 = x^2 + 9x +18 0 = [x^2 + 3x] + [6x +18] 0 =x [x + 3] + 6[x +3] 0 =[x + 3] [x +6] Hence x = -3 or x = -6

OpenStudy (anonymous):

i got x=-6

OpenStudy (anonymous):

Yes but there are two roots